# Synomaly Noise and Multi-Stage Diffusion: A Novel Approach for Unsupervised Anomaly Detection in Ultrasound Imaging

### v1

@@ 4.1 Datasets
In this work, three datasets are used, namely, a brain MRI dataset, a liver CT dataset, and a carotid US dataset. The details of each dataset are provided below.
4.1.1 Brain MRI: BraTS Dataset
The BraTS2023 dataset originates from the international Brain Tumor Segmentation (BraTS) Challenge [2, 3, 31], which is a well-established benchmark in the field of medical imaging. It comprises 3D brain MRI scans from patients diagnosed with brain tumors, providing a rich and diverse set of data for developing and evaluating segmentation models. The pixel-wise annotations classify tissues into the four categories healthy tissue, GD-enhancing tumor (ET), peritumoral edematous/invaded tissue (ED) and necrotic tumor core (NCR). For the purposes of this work, all three tumor-related classes (ET, ED, and NCR) are grouped into a single “anomalous” category, forming two classes: healthy and anomalous tissue.
Since the focus is on tumor detection in 2D images, only axial brain slices are considered in this work and the 2D tumor ground truth segmentation is extracted from the annotated 3D classification volumes. Given that ET, ED and NCR are most commonly found in the cerebral hemispheres, the brain MRI dataset is further refined to include only the central slices. For anomaly detection, the FLAIR sequence is used. All selected slices are resized to pixels to have unified dimensions. anomalous slices are randomly selected as the test set. In total, the dataset consists of healthy training samples and anomalous test samples. To help intuitive understanding, a few representative samples are visualized in Fig. 4(b).
4.1.2 Liver CT: LiTS Dataset
The Liver Tumor Segmentation Benchmark (LiTS) dataset [9] includes abdominal CT volumes, each with detailed annotations of the liver and its tumors. The number of tumors per volume ranges from to , containing both primary and secondary tumors of diverse sizes and appearances. For the purposes of this work, all tumors are categorized as “anomalous” class.
Each 3D abdominal CT volume contains varying axial slices, ranging from and . To prepare the data for model training and evaluation, the annotated segmentations are used to classify slices into healthy and anomalous cases. Further refinement is performed using size thresholding to isolate the relevant slices and all images are resized to pixels. In total, the CT dataset has healthy training images and anomalous test images. A set of representative samples is shown in Fig. 4(c). These examples highlight the complexity of the anomaly detection task, as the abdominal region exhibits various appearances and the liver itself can differ in size, shape and location. The relatively small ratio between the anomaly and the chest volume can further cause class imbalance problems. These factors make it particularly challenging to accurately segment liver tumors, while preserving healthy tissue. This high diversity and complexity of the LiTS dataset make it a great benchmark for assessing the performance and generalizability of the proposed anomaly detection approaches.
4.1.3 Carotid US Dataset
The carotid artery US dataset includes US scans from healthy patients and patients with carotid plaques. The US images in the dataset are all captured from the cross-sectional view of the carotid artery. The acquisition was performed within the Institutional Review Board Approval by the Ethical Commission of the Technical University of Munich (reference number 244/19 S). Due to the nature of US imaging, B-mode appearance is affected significantly by contact point, force, angle, and US machine setting. So, to alleviate the effects caused by the changing image appearance, we crop the US image around the target anatomy, i.e., the carotid artery, to ensure the network only focuses on disease-cased image variations. In this study, we applied existing vessel tracking and segmentation methods to generate vessel masks, which are then used to crop the original full-view US images. Fig. 4(a) presents the pipeline for vessel tracking and cropping procedure. To account for potential imperfections in vessel segmentation and to enhance model robustness, a small degree of randomness regarding the position and size of the vessel within the image is introduced in the cropping process, which also creates some variability in the training data.
The ground truth annotations for plaques in the anomalous carotid artery samples are manually annotated under the supervision of our clinical experts specialized in vascular disease. All such annotations of vessels and anomalies were performed using the ImFusion Suite111https://www.imfusion.com/products/imfusion-suite. All US images are resized to pixels to standardize input dimensions. In total, the final dataset contains healthy training images and anomalous test images. The examples can be found in Fig. 4(d).
@@ 4.2 Inference Hyperparameters
In summary, the inference noise steps for partial diffusion (), the Gaussian kernel size for blurring (), and the anomaly threshold for anomaly selection () are the three hyperparameters that need to be carefully tuned for each dataset and diffusion noise function. To ensure a fair comparison between different models, the best inference hyperparameter is found for each model, representing the upper limit of performance.
@@ 4.3 Training Details
The performance of the proposed anomaly detection method is compared with six state-of-the-art (SOTA) methods: classic VAE, DDPM using Gaussian noise function [23], DDPM using Coarse noise function [26], DDPM using Pyramid noise function [20], DDPM using Simplex noise function [42], AutoDDPM [7], and DAE [25]. Fully supervised UNets [37] were also trained for each dataset as a reference. Half of the anomalous data of each dataset is selected as UNet training data.
@@
The VAE model and DAE model are trained for epochs using Adam optimizer with an exponentially decreasing learning rate from to by . All the DDPM models are implemented based on the popular guided diffusion approach by OpenAI [12]. For the Simplex noise function, the noising parameters were set with an octave of 6, a persistence of 0.8 and a frequency of 64, as recommended by the original authors. For Coarse noise, the noising parameters were configured with a noise resolution of 16 and a noise standard deviation of 0.2, also aligning with the established literature. All the DDPMs are trained using Adam optimizer with a fixed learning rate of . The models were trained for 3000 epochs on the brain MRI and liver CT datasets, and for 2000 epochs on the carotid US dataset. All the models were trained using Nvidia A40 GPU.
4.4 Results on Brain MRI
@@
For the anomaly detection in brain MRI images, the Synomaly noise function parameters are configured with a blurring variance of 3, a threshold value of 175, an anomaly direction of 1 and an intensity offset of 0.5. These Synomaly parameters were determined through a combination of known disease morphology and experimental validation. To ensure a fair comparison between methods, the optimal inference hyperparameters for each method were determined through a comprehensive grid search. Tab. 1 presents the best hyperparameter combinations for each method. Notably, for VAE and DAE, since no diffusion process is involved, the “noise steps” field is left blank. “DDPM+Synomaly+Multi-stage” depicts the case where the proposed Synomaly noise is used to train the DDPM and the multi-stage diffusion process is utilized during inference. The “noise steps” field for “DDPM+Synomaly+Multi-stage” indicates the noise steps at each stage of the iterative process.
@@
Using the best parameters identified, each model was tested on 1000 anomalous images. The evaluation metrics include the Dice, Precision, and Recall score, to ensure an objective and comprehensive comparison of different approaches. The anomaly segmentation results are shown in Tab. 2. It is evident that the proposed Synomaly noise function consistently outperforms all other unsupervised methods across all metrics. Fig. 5 shows the visual results of the reconstructed images and anomaly segmentation maps using different unsupervised methods. DDPMs trained using Coarse noise and Simplex noise functions partially erase the anomalous regions, but distinct borders of the brain tumors remain in the reconstructed images. For AutoDDPM and DAE, the bright anomalous area is slightly dimmed but not completely removed. However, changes in these areas can still be identified with a proper threshold. The VAE reconstruction is anomaly-free but appears blurry, losing important brain details. Meanwhile, DDPMs trained with Gaussian noise and Pyramid noise fully remove the round anomaly in the top area of the image but fail to accurately reconstruct the lateral ventricle. In contrast, the image reconstructed using the proposed method shows a more realistic reconstruction, particularly for the lateral ventricle, demonstrating its superior performance in terms of generative fidelity. The proposed approach delivers the best anomaly segmentation mask compared to the other SOTAs.
4.5 Results on Liver CT
@@
The configuration of Synomaly noise function for liver CT anomaly detection task is set as a blurring variance of 3, a threshold value of 175, an anomaly direction of -1 and an intensity offset of 1. Note that for this specific use case, the anomaly detection direction was set to -1 instead of 1, since the liver tumors appear darker than the surrounding healthy tissue, requiring an inversion of the detection parameter.
@@
The best inference hyperparameters are determined for each method by applying grid search. The anomaly segmentation performance is summarized in Tab. 2. In this scenario, the VAE completely fails to generate plausible counterfactual samples, as the complexity of CT images exceeds the generative capacity of the VAE model. As shown in Fig. 5, the VAE-generated images are of low resolution with jagged borders. The DDPMs trained with Gaussian and Pyramid noise function can effectively erase the abnormalities but they also corrupt other parts of the image, resulting in a high false-positive detection rate. The AutoDDPM and DAE cannot fully remove the anomalies, but with a proper threshold, the anomalous regions can still be highlighted. The DDPM trained with the Simplex noise function demonstrated relatively good generation results, with liver tumors almost fully removed. The best performance was observed with Synomaly noise plus the multi-stage diffusion process, which achieved the highest scores across all metrics. The visual results also confirm its superior performance, as the anomalies are completely erased without any trace. When comparing the performance of unsupervised anomaly segmentation methods to fully supervised segmentation models, a significant gap remains in this scenario. This can be attributed to the complexity of abdominal CT images, where the generative model must reconstruct not only the liver but also all other organs in the abdomen, making accurate segmentation considerably more challenging.
4.6 Results on Carotid US
@@
The Synomaly noise function in this scenario was defined with parameters based on known disease morphology and experimental validation, which resulted in a blurring variance of 7, a threshold value of 150, an anomaly direction of 1 and an intensity offset of 0.5. To identify the best inference parameters for each model, a comprehensive grid search was conducted across the parameters noise steps (), Gaussian kernel size () and anomaly threshold (). Tab. 1 presents the optimal parameters for each method.
@@
The anomaly segmentation results are summarized in Tab. 2. It is evident that Synomaly noise significantly outperforms all other unsupervised methods. Interestingly, the best performance apart from the proposed approach in this scenario is achieved by the VAE, with an average Dice score of . This can be attributed to the fact that, compared to brain MRI and liver CT, the structure of the carotid vessel is relatively simple, allowing the VAE to reconstruct a plausible plaque-free artery from a diseased one. Additionally, carotid plaques are more distinct and larger in size, which makes removing such anomalies with traditional SOTA noise functions challenging. Large noising steps are required, resulting in severe corruption of the image and significant information loss. Due to the similarity in textures between plaques and surrounding tissues, noising further blurs the distinction between these structures. As observed in the fifth row of Fig. 5, generated vessels often incorrectly interpret the inner border of the original diseased vessel as the outer border, leading to reconstructions of smaller-than-actual arteries. The DAE fails to effectively remove the plaques partly because the added noise during training does not sufficiently represent the anomalies. Consequently, during inference, the network encounters situations that it was not exposed to during training. In contrast, Synomaly noise effectively identifies and removes the plaques within the artery. The image generated using the proposed approach demonstrates the best performance among all methods evaluated. As illustrated in the last subplot of the fifth row of Fig. 5, the plaques are completely removed, and the artery’s borders are accurately reconstructed.
@@ 4.7 Ablation Study
In this section, we conduct ablation studies on the hyperparameters of Synomaly noise and the multi-stage diffusion process. These experiments aim to provide readers with a clearer understanding of how these hyperparameters influence the final anomaly segmentation performance and highlight the importance of employing the multi-stage diffusion process.
4.7.1 Hyperparameter Ablation for Synomaly Noise
In this ablation study, we first examine the influence of the hyperparameters of Synomaly noise, as described in Sec. 3.2. The performance of Synomaly noise is closely linked to its ability to represent the target anomalies. If the Synomaly noise generated during training resembles the real abnormalities encountered during inference, the trained model can more effectively erase the diseased areas during testing. The morphology of the generated Synomaly noise is controlled by two parameters: the blurring variance () and the threshold value ().
To evaluate the influence of these parameters, several models were trained with different values of and on the carotid US dataset. The generated noise is normalized to the range [0, 255], and consequently, the threshold values are also within this range. First, we kept fixed at 150 and varied from 1 to 11. The results are presented in Fig. 6(a), which shows that the best performance was achieved with a blurring variance of 7. Similarly, for , we conducted experiments by fixing at 7 while varying from 90 to 230. The results, as shown in Fig. 6(b), indicate that the best performance was achieved when setting to 150. Some examples of the generated Synomaly noise are displayed below the plots. The value of controls the size of the synthetic anomaly regions, while determines the number of generated artificial anomaly areas. A smaller value results in smaller synthetic anomaly regions, whereas a lower value leads to a higher number of generated anomaly areas. However, in some cases, when choosing a smaller value, these generated regions sometimes overlap with each other, leading to fewer but larger anomalous areas. For the case of US plaque detection, the optimal choice is , , since this combination can best resemble the carotid plaque morphology based on our experiments.

### v2

@@ 4.1 Datasets
We evaluate the proposed method on three representative datasets spanning different modalities and anatomical regions: Brain MRI (BraTS 2023) [2, 3, 44], Liver CT (LiTS) [13], and Carotid US. The datasets comprise both healthy and anomalous cases, with Brain MRI containing 3743 healthy training slices, 1000 anomalous test slices, and 416 healthy test slices; Liver CT including 5820 healthy training images, 1000 anomalous test images, and 647 healthy test images; and Carotid US consisting of 7306 healthy training images, 545 anomalous test images, and 833 healthy test images. Additional details on data acquisition, preprocessing, and sample images are provided in Appendix A.
@@ 4.2 Inference Hyperparameters
In summary, the inference noise steps for partial diffusion (), the Gaussian kernel size for blurring (), and the anomaly threshold for anomaly selection () are the three primary hyperparameters considered during inference. However, it is notable that the Gaussian kernel size and threshold value do not impact the quality of the generated counterfactual image. These parameters are used solely in the post-processing step to refine the anomaly mask by controlling the level of smoothing and the sensitivity of anomaly detection. In contrast, the number of noise steps plays a crucial role in determining the balance between preserving fine details and effectively removing anomalies during counterfactual image reconstruction. To ensure a fair comparison between different models, we determine the optimal inference hyperparameters for each approach, representing their best possible performance. As shown in Tab. 1, the proposed method (DDPM+Synomaly+Multi-stage) shows minimal variation in noise steps across datasets, differing by only 100 steps, demonstrating its robustness to hyperparameter tuning.
@@ 4.3 Training Details
The performance of the proposed anomaly detection method is compared with eight state-of-the-art (SOTA) methods: classic VAE [36], f-AnoGAN [52], NSA [54], DDPM using Gaussian noise function [26], DDPM using Coarse noise function [34], DDPM using Pyramid noise function [24], DDPM using Simplex noise function [61] (also known as AnoDDPM), AutoDDPM [8], and DAE [33]. Fully supervised UNets [51] were also trained for each dataset as a reference. Half of the anomalous data of each dataset is selected as UNet training data.
| Method | Brain MRI BraTS23 | Liver CT LiTS | Carotid Plaque US | ||||||
| DSC | Precision | Recall | DSC | Precision | Recall | DSC | Precision | Recall | |
| VAE [51] | |||||||||
| f-AnoGAN [52] | |||||||||
| NSA [54] | |||||||||
| DDPM+Gaussian [26] | |||||||||
| DDPM+Corase [34] | |||||||||
| DDPM+Pyramid [24] | |||||||||
| DDPM+Simplex [61] | |||||||||
| AutoDDPM [8] | |||||||||
| DAE [33] | |||||||||
| DDPM+Synomaly+Multi-stage | |||||||||
| UNet [51]∗ | |||||||||
Fully supervised method.
@@
The VAE model, f-AnoGAN, NSA model, and DAE model are trained for epochs using Adam optimizer with an exponentially decreasing learning rate from to by . All the DDPM models are implemented based on the popular guided diffusion approach by OpenAI [17]. For the Simplex noise function, the noising parameters were set with an octave of 6, a persistence of 0.8 and a frequency of 64, as recommended by the original authors. For Coarse noise, the noising parameters were configured with a noise resolution of 16 and a noise standard deviation of 0.2, also aligning with the established literature. All the DDPMs are trained using Adam optimizer with a fixed learning rate of . The models were trained for 3000 epochs on the brain MRI and liver CT datasets, and for 2000 epochs on the carotid US dataset. All the models were trained using Nvidia A40 GPU. During inference, DDIM [56] is applied to accelerate the inference process.
4.4 Pixel-level Anomaly Detection Performance
In this section, we evaluate the performance of the proposed method on pixel-level anomaly detection, comparing it with SOTA methods. The evaluation is conducted using anomalous images from the test set, where segmentation performance is measured to assess how effectively each method detects and localizes anomalies. Quantitative results are reported in Tab. 2, while the visual results are shown in Fig. 4. Since NSA is not a generative model and only outputs anomaly masks, its results are not shown in Fig. 4.
4.4.1 Results on Brain MRI
@@
For the anomaly detection in brain MRI images, the Synomaly noise function parameters are configured with a blurring variance of 3, a threshold value of 175, an anomaly direction of 1 and an intensity offset of 0.5. These Synomaly parameters were determined through a combination of known disease morphology and experimental validation. No mask is applied when generating Synomaly noise for brain MRI, as the brain occupies the majority of the image and is easily distinguishable from the background, reducing the risk of anomalies being misplaced in irrelevant areas. To ensure a fair comparison between methods, the optimal inference hyperparameters for each method were determined through a comprehensive grid search. Tab. 1 presents the best hyperparameter combinations for each method. Notably, for VAE and DAE, since no diffusion process is involved, the “noise steps” field is left blank. “DDPM+Synomaly+Multi-stage” depicts the case where the proposed Synomaly noise is used to train the DDPM and the multi-stage diffusion process is utilized during inference. The “noise steps” field for “DDPM+Synomaly+Multi-stage” indicates the noise steps at each stage of the iterative process.
@@
Using the best parameters identified, each model was tested on 1000 anomalous images. The evaluation metrics include the Dice, Precision, and Recall score, to ensure an objective and comprehensive comparison of different approaches. The anomaly segmentation results are shown in Tab. 2. It is evident that the proposed method consistently achieves the highest performance in terms of DSC and precision. While NSA yields the highest recall in the brain MRI dataset, its substantially lower precision suggests that it tends to over-segment, reducing its overall effectiveness. Fig. 4 shows the visual results of the reconstructed images and anomaly segmentation maps using different unsupervised methods. DDPMs trained using Coarse noise and Simplex noise functions partially erase the anomalous regions, but distinct borders of the brain tumors remain in the reconstructed images. For AutoDDPM and DAE, the bright anomalous area is slightly dimmed but not completely removed. However, changes in these areas can still be identified with a proper threshold. The reconstruction results of VAE and f-AnoGAN are anomaly-free but appears blurry, losing important brain details. Meanwhile, DDPMs trained with Gaussian noise and Pyramid noise fully remove the round anomaly in the top area of the image but fail to accurately reconstruct the lateral ventricle. In contrast, the image reconstructed using the proposed method shows a more realistic reconstruction, particularly for the lateral ventricle, demonstrating its superior performance in terms of generative fidelity. The proposed approach delivers the best anomaly segmentation mask compared to the other SOTAs.
4.4.2 Results on Liver CT
@@
The configuration of Synomaly noise function for liver CT anomaly detection task is set as a blurring variance of 3, a threshold value of 175, an anomaly direction of -1 and an intensity offset of 1. Note that for this specific use case, the anomaly detection direction was set to -1 instead of 1, since the liver tumors appear darker than the surrounding healthy tissue, requiring an inversion of the detection parameter. The generated Synomaly noise is masked using the liver segmentation provided in the dataset, ensuring that synthetic anomalies are predominantly placed inside the liver region.
@@
The best inference hyperparameters are determined for each method by applying grid search. The anomaly segmentation performance is summarized in Tab. 2. In this scenario, the VAE, f-AnoGAN, and NSA completely fail to generate plausible counterfactual samples, as the complexity of CT images exceeds the generative capacity of the VAE model. As shown in Fig. 4, the generated images of VAE and f-AnoGAN are of low resolution with jagged borders. The DDPMs trained with Gaussian and Pyramid noise function can effectively erase the abnormalities but they also corrupt other parts of the image, resulting in a high false-positive detection rate. The AutoDDPM and DAE cannot fully remove the anomalies, but with a proper threshold, the anomalous regions can still be highlighted. The DDPM trained with the Simplex noise function demonstrated relatively good generation results, with liver tumors almost fully removed. The best performance was observed with Synomaly noise plus the multi-stage diffusion process, which achieved the highest scores in DSC and precision. The visual results also confirm its superior performance, as the anomalies are completely erased without any trace. When comparing the performance of unsupervised anomaly segmentation methods to fully supervised segmentation models, a significant gap remains in this scenario. This can be attributed to the complexity of abdominal CT images, where the generative model must reconstruct not only the liver but also all other organs in the abdomen, making accurate segmentation considerably more challenging.
4.4.3 Results on Carotid US
@@
The Synomaly noise function in this scenario was defined with parameters based on known disease morphology and experimental validation, which resulted in a blurring variance of 7, a threshold value of 150, an anomaly direction of 1 and an intensity offset of 0.5. A circular mask is applied to restrict synthetic anomalies within the vessel region, ensuring that the generated anomalies align with the expected anatomical distribution. The mask has a diameter of 90% of the image size and is centered within the image to approximate the vessel’s location. To identify the best inference parameters for each model, a comprehensive grid search was conducted across the parameters noise steps (), Gaussian kernel size () and anomaly threshold (). Tab. 1 presents the optimal parameters for each method.
@@
The anomaly segmentation results are summarized in Tab. 2. It is evident that Synomaly noise significantly outperforms all other unsupervised methods. Interestingly, the best performance apart from the proposed approach in this scenario is achieved by the VAE, with an average Dice score of . This can be attributed to the fact that, compared to brain MRI and liver CT, the structure of the carotid vessel is relatively simple, allowing the generative models like GANs and VAEs to reconstruct a plausible plaque-free artery from a diseased one. Additionally, carotid plaques are more distinct and larger in size, which makes removing such anomalies with traditional SOTA noise functions challenging. Large noising steps are required, resulting in severe corruption of the image and significant information loss. Due to the similarity in textures between plaques and surrounding tissues, noising further blurs the distinction between these structures. As observed in the fifth row of Fig. 4, generated vessels often incorrectly interpret the inner border of the original diseased vessel as the outer border, leading to reconstructions of smaller-than-actual arteries. The DAE fails to effectively remove the plaques partly because the added noise during training does not sufficiently represent the anomalies. Consequently, during inference, the network encounters situations that it was not exposed to during training. In contrast, Synomaly noise effectively identifies and removes the plaques within the artery. The image generated using the proposed approach demonstrates the best performance among all methods evaluated. As illustrated in the last subplot of the fifth row of Fig. 4, the plaques are completely removed, and the artery’s borders are accurately reconstructed.
@@
We acknowledge that, to reduce complexity, a vessel tracking step was applied to localize the carotid artery region and minimize irrelevant background information. This preprocessing mitigates the high variability of US backgrounds, which can mislead generative models. Unlike CT and MRI, where background structures provide a stable anatomical reference, US backgrounds are inconsistent, making anomaly detection more challenging. Despite this simplification, the task remains difficult—even a fully supervised segmentation model trained on zoomed-in anomalous images achieved only 0.797 Dice, highlighting that anomaly detection remains a challenging task even after vessel localization.
To demonstrate that vessel tracking can be achieved with minimal effort, we conducted an additional experiment using Segment Anything Model 2 (SAM2) [50] for automated carotid vessel tracking before applying the proposed anomaly detection method. Only the ground truth vessel mask of the first frame in each US sweep was provided to SAM2 for initialization, after which the model tracked the vessel across subsequent frames, and images were cropped based on the tracking results. Without any fine-tuning on US data, SAM2 successfully tracked the carotid region. Based on the tracking results, the final anomaly segmentation achieved a Dice score of . This suggests that modern large segmentation models can perform robust vessel tracking with minimal manual intervention. However, while vessel tracking benefits from strong shape priors, this may not generalize to more irregular structures like anomalies, where tracking remains challenging. These results highlight the feasibility of automated tracking as a practical preprocessing step for improving anomaly detection in US imaging.
4.5 Image-level Anomaly Detection Performance
In this section, we evaluate the performance of the proposed method on image-level anomaly detection, assessing its ability to distinguish anomalous images from healthy ones. The evaluation is conducted using both healthy and anomalous images in the test set, and performance is measured using the Area Under the Receiver Operating Characteristic Curve (AUROC). The results are compared across different imaging modalities to demonstrate the method’s effectiveness in various clinical scenarios.
The AUROC scores achieved by the proposed method are 0.922 for brain MRI, 0.920 for liver CT, and 0.991 for carotid US, as illustrated in Fig. 5. In the figure, the blue curve represents the AUROC for brain MRI, the orange curve corresponds to liver CT, and the green curve represents carotid US. These results indicate that the proposed method consistently performs well across different imaging modalities in distinguishing anomalous from healthy images. The particularly high AUROC score for carotid US (0.991) suggests that the method is highly effective in this modality, likely due to the more challenging appearance of anomalies compared to normal structures.
While these results confirm the strong classification capability of the proposed method, pixel-level anomaly segmentation provides additional benefits by offering localized information about abnormal regions, which can improve interpretability and aid in detailed analysis. Additionally, pixel-level AD is inherently more challenging than image-level classification, as it requires the model to distinguish fine-grained abnormal structures from normal variations within an image. Nonetheless, the inclusion of image-level classification metrics further supports the overall effectiveness of the approach and provides a broader evaluation perspective.
@@ 4.6 Ablation Study
In this section, we conduct ablation studies to evaluate the impact of three key components of our method: the multi-stage diffusion process, Synomaly noise, and the hyperparameter selection for Synomaly noise. First, we analyze the effect of the multi-stage diffusion process by comparing it to the single-stage approach, assessing its role in improving anomaly removal while preserving fine details. Next, we evaluate the contribution of Synomaly noise to anomaly segmentation by comparing it with alternative noise functions. Finally, we examine the impact of hyperparameter selection for Synomaly noise, investigating how different configurations influence segmentation performance. These studies provide a deeper understanding of how each component contributes to the overall effectiveness of the proposed method.